option parsing all options passed to program loading input files input files: genomic sequence reference sequence splice site param. similarity filter List of Hits DP parameter  DP algorithm for optimal spliced alignment  one spliced alignment spliced alignment collection binary search tree of spliced alignments output spliced alignments output of spliced alignments BSSM parameter files  mkbssmfiles.x index of genomic and reference sequence mkvtree.x  for every hit  possibly